OpenStudy (anonymous):

A truck needs 7 gallons of fuel to travel 56 miles. Can the truck travel 48 miles with 6 gallons of fuel? Explain. a.Yes;7/56 and 6/48are proportional because 7*48=6*56 b.Yes; 7/56 and48/6 are proportional because 7*6 = 48*56 c.No;7/56 and 6/48 are not proportional because d.No

OpenStudy (anonymous):

a.Yes;7/56 and 6/48are proportional because 7*48=6*56

OpenStudy (anonymous):

c.\[7\times48\neq6\times56\] d.\[7\times6\neq48\times56\]

OpenStudy (anonymous):

c. 7×48≠6×56

OpenStudy (anonymous):

A 58/7=8 This means the truck gets eight miles to the gallon... so eight times six is 48 exactly... The truck would be able to go exactly 48 miles on six gallons of gas The answeris A

OpenStudy (anonymous):

The problem was solved...

OpenStudy (anonymous):

Find a number that is a solution of the inequality \[ x=\le1\]

OpenStudy (anonymous):

@SpartanRace